OpenStudy (mackenzie2013):

Eight two-person teams are participating in a three-legged race. Find the number of orders in which all 8 teams can finish the race.

OpenStudy (anonymous):

\[8!\]

OpenStudy (anonymous):

8 choices to come in first 7 for secone 6 for third 5 for fourth 4 for fifth 3 for sixth 2 for seventh then 1 for eighth by the counting principle, multiply the choices together, get \[8!=8\times 7\times 6\times 5\times 4\times 3\times 2\times 1\]

OpenStudy (mackenzie2013):

Thank you!